exercise 2you decide to sell short 100 shares of Charlotte Horse Farms when it is selling at its yearly high of $56

exercise 2
you decide to sell short 100 shares of Charlotte Horse Farms when it is selling at its yearly high of $56 . Your broker tells you that your margin requirement is 45 percent and the commission on the purchase is $155 . While you are short the stock , Charlotte pays a $2.50 per share dividend. At the end of the one year, you buy 100 shares of Charlotte at $45 to close out your position and are charged a commission of $145 and 8 percent interest on the money borrowed .What is your rate of return on the investment.

Exercise 2
You own 200 shares of Shamrock Enterprises that you bought at $25 a share. The stock is now selling for $45 a share.
• You put in a stop loss order at $40 .Discuss your reasons for this action.
• If the stock eventually declines in price to $30 a share, what would be your rate of return with and without the stop loss order ?
Answer #1

BEGIBIBG VALUE=56 X100 =5600

INITIAL INVESTMENT =MARGINAL REQUIREMENT +COMMISION

= .45 X5600 + 155= 2675

ENDING VALUE= 45 X 100=4500

DIIVIDEND =2.50 X100= 250

TRANSZCTIONS COSTS=155 +145 =300

INTEREST =0.8 X (.55 X 5600)= 246 .40

PROFIT =5600 -4500-250-246.40= 303.60

RATE OF RETURN ON YOUR INVESTMENT =303.6/ 2675=11.35%
Answer #2
BEGIBIBG VALUE=56 X100 =5600 INITIAL INVESTMENT =MARGINAL REQUIREMENT
+COMMISION = .45 X5600 + 155= 2675 ENDING VALUE= 45 X 100=4500 DIIVIDEND =2.50 X100= 250 TRANSZCTIONS COSTS=155 +145 =300 INTEREST =0.8 X (.55 X 5600)= 246 .40 PROFIT =5600 -4500-250-246.40= 303.60 RATE OF RETURN ON YOUR INVESTMENT =303.6/
2675=11.35%
Answer #3
Answer Details
Interest on money borrowed = 100*56*45%*8% = 201.6
Commission = $155

Dividends = 100 * 2.5 = 250
Gain at the year end on short sale of stock = 100*(56-45) = 1100
Commsion = 145

Net gain = 1100-201.6-155+250-145 = 848.4 Answer

Happy to hel;p